The permiter of the garden shown is 6x+2 units. Find the missing length of the missing side.
gives me on side x one side x and bottom 3x+1 missing the top length

Answer by richwmiller(17219)

6x+2-x-x-(3x+1)=s
4x+2-3x-1=s
x+1=s